### v1

# Benchmarking Multi-Modal Retrieval for Long Documents

Abstract.
Multi-modal document retrieval is designed to identify and retrieve various forms of multi-modal content, such as figures, tables, charts, and layout information from extensive documents. Despite its significance, there is a notable lack of a robust benchmark to effectively evaluate the performance of systems in multi-modal document retrieval. To address this gap, this work introduces a new benchmark, named as MMDocIR, encompassing two distinct tasks: page-level and layout-level retrieval. The former focuses on localizing the most relevant pages within a long document, while the latter targets the detection of specific layouts, offering a more fine-grained granularity than whole-page analysis. A layout can refer to a variety of elements such as textual paragraphs, equations, figures, tables, or charts. The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation. Through rigorous experiments, we reveal that (i) visual retrievers significantly outperform their text counterparts, (ii) MMDocIR train set can effectively benefit the training process of multi-modal document retrieval and (iii) text retrievers leveraging on VLM-text perform much better than those using OCR-text. These findings underscores the potential advantages of integrating visual elements for multi-modal document retrieval. MMDocIR is available at https://huggingface.co/MMDocIR.
Multi-modal Document Retrieval, Page Retrieval, Layout Retrieval
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Multi-modal Document Retrieval
Figure 1. MMDocIR comprises 313 lengthy documents across 10 different domains, along with 1,685 questions. For each question, page-level annotations are provided via selected screenshots. Red boundary boxes represent layout-level annotations.
Figure 2. Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark (Ma et al., 2024b). Note that the white spaces, headers, and footers are removed from the area counting.
1. Introduction
Multi-modal document retrieval (Hassan et al., 2013; Ma et al., 2024a; Faysse et al., 2024; Lee et al., 2024) aims to encode and retrieve various forms of content from visually rich documents based on user queries. Unlike traditional document retrieval (Zhang et al., 2022; Chen et al., 2023; Sarthi et al., 2024; Dong et al., 2024) which primarily deals with textual data, multi-modal document retrieval must handle documents that seamlessly integrate text with multi-modal elements such as images, tables, charts, and layout designs. These multi-modal content often carry significant information that plain text cannot convey (Cui et al., 2021; Zhu et al., 2024; Sassioui et al., 2023), thereby enriching the depth and relevance of retrieved content. For example, our analysis of MMLongBench-Doc benchmark (Ma et al., 2024b) reveals that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively, as shown in Figure 2. This significant presence of multiple data modalities requires retriever to leverage all data modalities present in a document.
However, current benchmarks (shown in Table 1) for evaluating multi-modal document retrievers are insufficient, lacking in certain aspects that are critical for a comprehensive assessment. The major shortcomings include: 1. Question Quality: The design and curation of questions in most benchmarks do not align with the specific needs of multi-modal document retrieval. The questions, often reused from Visual Question Answering (VQA) tasks, lack the specificity required for this purpose and require expert filtering to be truly effective for retrieval tasks. 2. Document Quality: A significant limitation of existing benchmarks is the missing of complete document pages, thus hindering accurate evaluation of retrieval capabilities within actual document contexts. Additionally, the range of document domains covered is often too narrow, limiting the benchmarks’ applicability across different fields. 3. Retrieval Granularity: Most benchmarks, except for SciMMIR (Wu et al., 2024), only allow retrieval at a page level which is insufficient, as user queries may pertain to specific parts of a page like a particular figure or table rather than the entire page.
Recognizing these gaps, we introduce MMDocIR Multi-Modal Document Information Retrieval benchmark. MMDocIR is innovatively structured around two critical tasks: page-level and layout-level retrieval. (1) The page-level retrieval task is designed to identify the most relevant pages within a document in response to a user query. (2) The layout-level retrieval aims to retrieve most relevant layouts. The layouts are defined as the fine-grained elements such as paragraphs, equations, figures, tables, and charts. This task allows for a more nuanced content retrieval, honing in on specific information that directly answers user queries. To support these tasks, we first develop MMDocIR evaluation set comprising 313 documents averaging 65.1 pages each, along with VQA 1,658 questions derived from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024). Initially, we examine all VQA questions, filter out questions that are irrelevant to IR tasks, and revise remaining questions to be more suitable for IR. Then, we manually annotate these questions with page and layout labels. The page-level labels are the specific pages identified that contain answer evidence.  [^1] It is worth noting that while MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in about 21.3% of these labels. The layout-level labels are the bounding boxes meticulously delineated around the key evidence within the identified pages. Meanwhile, we introduce MMDocIR training set, comprising 73,843 QA pairs converted from 7 DocQA datasets. Specifically, we manually collect 6,878 documents and use either manual or automatic methods to annotate the ground truth labels.  [^2] Since layout annotations are missing for most datasets, our data annotating pipeline can only create layout-level labels for 4 subsets.
Benchmarks Question Document Label
Type By Expert? For IR? #Num Evidence Type Domain #Pages Source Page Layout
DocCVQA (Tito et al., 2021) VQA question 20 TXT/L Finance 1.0
SciMMIR (Wu et al., 2024) Image caption 530K TAB/I Science 1.0
ViDoRe (Faysse et al., 2024) VQA question 3,810 TXT/C/TAB/I Multi-domain 1.0
PDF-MVQA (Ding et al., 2024) Search query 260k TXT/TAB/I Biomedical 9.6
MMLongBench-Doc (Ma et al., 2024b) VQA question 1,082 TXT/L/C/TAB/I Multi-domain 47.5
Wiki-SS (Ma et al., 2024a) Natural question 3,610 TXT Wikipedia 1.0
DocMatix-IR (Ma et al., 2024a) VQA question 5.61M TXT/L/C/TAB/I Multi-domain 4.2
MMDocIR VQA question 1,658 TXT/C/TAB/I Multi-domain 65.1
Table 1. Comparison between our benchmark and previous Document IR datasets. TXT/L/C/TAB/I: pure text/generalized layout/chart/table/image.
In our research, we conduct a thorough evaluation of the prevailing multi-modal document retrieval baselines which can broadly be categorized into visual-driven and text-driven retrievers. Visual-driven retrievers (Ma et al., 2024a; Faysse et al., 2024), leverage vision-language models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) to capture rich multi-modal cues and generate embeddings for both queries and documents. In contrast, text-driven retrievers (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Izacard et al., 2022; Wang et al., 2022; Xiao et al., 2023; Li et al., 2023) primarily rely on OCR or VLM to first convert the multi-modal content into text, subsequently employing language models (LMs)  (Devlin et al., 2019; Liu et al., 2019; Wang et al., 2024b) to generate embeddings for both queries and documents. Via extensive experiments, we discover that visual-driven retrievers consistently outperform their text-driven counterparts, often by a significant margin. This highlights the critical role that visual elements play in multi-modal document retrieval. In summary, our contributions are three-fold:
  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that encompasses page-level and fine-grained layout-level multi-modal document retrieval.
  • MMDocIR Benchmark: We introduce Multi-Modal Document Information Retrieval (MMDocIR) benchmark. The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 QA pairs. The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 QA pairs.
  • Comprehensive Evaluation: We conduct an extensive evaluation of existing retrieval systems (§$\S$ 6). Our results reveal that visual-driven retrievers consistently outperform their text-driven counterparts, highlighting the importance of utilizing visual information in multi-modal document retrieval.
2. Dual-Task Retrieval Definition
Let $\mathcal{D}$ be a document corpora consisting of a set of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection [^3] A document page usually comprises about 5 to 15 layouts, depending on its complexity.. The objective is to perform document page-level and layout-level retrieval. Specifically, the task involves retrieving a subset of these pages and layouts that are most relevant to a query $Q$, identified by the top $k$ entries where $k$ is significantly smaller than $n$ or $m$. The relevance of pages ($p$) and layouts ($l$) to $Q$ is quantified by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively. The retrieval system can be decomposed into (1) an offline indexing phase in which each page and layout from $\mathcal{P}$ and $\mathcal{L}$ is transformed into a vector representation, and (2) an online querying phase in which a query $Q$ is converted into a vector form, and the vectorized query is then compared against the indexed vectors using the similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
3. MMDocIR: Evaluation Set
{tabu}
l—c@ c@ c@ c@ c@ c@ c—c@   c@  c@   c@   —c@   c@  c@   c@
Dataset Domain Document Statistics   Questions (%)   Modality Distribution (%)
#Doc #QA #Page #Lay #Page #Lay %Lay Text Image Table Lay/ Text Image Table Title
Label Label /Doc /Page Meta
MMDocIR Eval 313 1,658 2,107 2,638 65.1 8.3 41.8 44.7 21.7 37.4 11.5 60.4 18.8 16.7 4.1
[…]
- Research Report 34 200 318 400 39.4 6.0 39.1 45.0 17.5 74.5 13.5 45.6 40.0 5.9 8.4
- Admin & Industry 10 59 82 113 16.8 9.1 45.1 78.0 20.3 13.5 13.5 70.1 11.7 14.9 3.2
- Tut & Workshop 17 102 165 225 57.5 4.1 43.8 37.2 61.7 24.5 9.8 28.0 57.3 6.3 8.3
- Academic Paper 75 386 473 571 19.5 10.1 48.4 28.8 25.7 50.0 10.4 74.6 12.8 11.1 1.5
- Brochure 15 76 121 178 30.3 9.7 41.1 60.5 52.6 18.4 36.8 33.3 50.8 8.5 7.0
- Financial Report 51 343 394 477 169.5 9.2 44.8 28.0 13.1 54.5 5.3 60.3 7.9 29.2 2.6
- Guidebook 22 112 168 223 78.4 10.0 33.6 51.8 54.4 26.8 17.8 63.7 20.0 12.1 4.1
- Government 44 111 116 132 68.9 6.9 45.4 69.37 2.7 0 7.6 88.2 3.7 5.7 2.4
- Laws 44 132 133 149 58.5 6.0 31.2 62.1 0 10.6 27.3 83.8 1.6 12.3 2.2
- News 1 137 137 170 50.0 73.6 72.3 70.1 1.5 0 28.5 48.5 39.8 0.0 11.6
Table 2. Detailed statistics for MMDocIR evaluation set. “#Lay/Page” is the averaging layouts per page, reflecting page’s layout complexity. “%Lay” refers to the area ratio of useful layouts (excluding white spaces, headers, and footers) over entire page.
3.1. Document Corpora Collection
To facilitate the development of MMDocIR, we leverage resources from existing Document Visual Question Answering (DocVQA) benchmarks (Mathew et al., 2021; Tito et al., 2023; Tanaka et al., 2023; Zhu et al., 2022; Li et al., 2024; Landeghem et al., 2023; Ma et al., 2024b; Zou et al., 2024). Although DocVQA benchmarks are not primarily designed for IR, they offer valuable document corpora and questions that can be adapted for our purposes. To select useful DocVQA datasets, we adhere to following criteria:
  • Document Source: The dataset must include accessible original documents or sources for these documents. We need to access and enrich them to support more complex retrieval tasks.
  • Diverse Domain/Modality: The document collections must (1) encompass diverse domains e.g., academia, finance, publishing, and research, and (2) embrace multiple modalities, such as text, figures, tables, charts, and layouts.
  • Long Document: We seek documents of sufficient length as longer texts pose more significant challenges. This criterion can evaluate models in handling complex and lengthy documents.
  • QA Diversity and Comprehensiveness: The questions included in the dataset should be diverse and challenging. For example, some questions should require reasoning across both text and visual tables/figures, while some are multi-hop questions that require synthesizing information from multiple pages.
After considering aforementioned criteria, we choose to leverage on the document corpora and questions from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024). MMLongBench-Doc is a long-context, multi-modal benchmark comprising 1,091 questions constructed upon 135 documents averaging 47.5 pages. Its corpora include domains of research report, administration and industry, tutorial and workshop, academic paper, brochure, financial report, and guidebook. DocBench focuses on long document understanding, consisting of 1,102 questions constructed upon 229 documents averaging 77.5 pages. The corpora span across domains: academia, finance, government, laws, and news. Both datasets provide expert-annotated questions, requiring content evidence of various modalities (e.g., text, tables, charts and images, as well as locations layouts) to answer. Consequently, we obtain a total of 364 documents and 2,193 questions for our subsequent annotation.
3.2. Annotation Process
Question Filtering and Revision
To ensure that the questions in MMDocIR are optimally suited for document retrieval tasks. To this end, we identify four specific types of questions that do not align well with the objectives of IR. By filtering out or modifying these questions, we ensure the integrity and relevance of MMDocIR.
  • Summarization Questions: These questions, such as “What does this book mainly illustrate?” require a comprehensive understanding of large sections of a document or even the entire document. The broad scope makes it hard to pinpoint specific content effectively, which contradicts the precise nature of IR tasks.
  • Overwhelm Statistical Questions: Questions that demand extensive data computation or collation, such as “How many words are there in total in the paper?” also fall outside our scope. These require a level of statistical analysis that goes beyond retrieval.
  • Online Search Questions: For example, queries like “What is the Google Scholar citation count of the main author of this article?” rely on information from external online resources. Since our focus is on retrieving information that is internally available within the documents, these questions are excluded.
  • Unanswerable Questions: These are designed to test if models generate answers based on non-existent information (model hallucinations). Since they do not facilitate the retrieval of factual document-based information, these questions are excluded.
Page-level annotation process
To address the challenge of accurately locating relevant information within extensive documents, we annotate page-level labels that pinpoint the exact pages containing evidence needed to answer specific questions. Considering that the average document in our dataset contains 65.1 pages, pinpointing the correct pages is similar to finding a needle in a haystack and requires meticulous effort. Our annotation process is:
  • For MMLongBench-Doc: we review and correct the answers and corresponding page labels for 794 questions. Specifically, we make corrections to 10 answers and revise 169 page labels [^4] Common errors in page labeling: annotators starting page indexing at 1, missing labels for questions spanning multiple pages, and incorrect or absent page labels. to ensure they accurately reflect the evidence locations.
  • For DocBench: we annotate page labels for all 864 questions from scratch. This involves a detailed examination of each document to accurately assign page labels that match the evidential requirements of each question.
Through these efforts and following a rigorous annotation and cross-validation process, we successfully compile page-level labels for a total of 1,658 questions in MMDocIR.
Layout-level annotation process
To achieve more fine-grained retrieval granularity, we go beyond page-level annotations by including layout-level labels. This advanced annotation process captures the specific layout elements within pages that contain necessary evidence to answer each question. Unlike page-level annotation, labeling layout-level information is more complex and labor-intensive due to its detailed nature. Our annotations are via:
  • To streamline annotation process, we utilize MinerU (Wang et al., 2024a) to parse all documents and extract the layout information (e.g., type and bounding boxes) for every layout element on a page. These elements are categorized into five main types: text, image, table, title, and equation. The bounding boxes are delineated by the coordinates of the top-right and bottom-left corners.
  • Based on the layouts drawn by MinerU, we identify the layouts that contain necessary eivdence to answer the question. In instances where MinerU fails to detect certain evidentiary elements, we manually draws the bounding boxes to ensure the precision of layout labels. This manual intervention is necessary for approximately 7% of our layout-level labels.
Ultimately, this meticulous process leads to the annotation of 2,638 layout labels for the 1,658 questions in MMDocIR.
3.3. Quality Control
Consistency Page Labels Layout Labels
Prec. Recall F1 Prec. Recall F1
[…]
Average 95.0 95.4 95.2 87.0 87.2 87.1
[…]
{tabu}
l—c—ccc@ c@ c—c@   c@  c@   c@   —c@   c@
Dataset Domain Document Statistics   Modality Distribution (%)   Labels
#Doc #QA #Page #Lay %Lay Text Image Table Title Page Lay
/Doc /Page
MMDocIR Train assorted docs 6,878 73,843 32.6 6.32 42.6 49.3 34.3 10.8 4.9
[…]
- MP-DocVQA (Tito et al., 2023) health/ind. docs 875 15,266 46.8 6.9 38.8 57.3 18.0 22.7 1.9
- SlideVQA (Tanaka et al., 2023) diverse slides 2,011 11,066 49.3 4.4 42.3 30.1 56.2 4.7 8.8
- TAT-DQA (Zhu et al., 2022) annual reports 163 15,814 147.3 9.2 42.2 66.4 4.4 26.5 2.7
- ArXivQA (Li et al., 2024) arXiv papers 1,579 12,314 18.4 7.9 50.0 70.4 22.3 2.8 1.0
- SciQAG (Wan et al., 2024) science papers 1,197 4,976 9.0 9.1 53.7 61.8 28.0 6.7 1.5
- DUDE (Landeghem et al., 2023) assorted docs 779 3,173 15.6 7.4 42.5 57.1 24.7 15.2 2.9
- CUAD (Hendrycks et al., 2021) legal contract 274 11,234 29.6 7.4 24.7 89.3 2.5 6.4 1.1
Table 4. Document statistics for Training Datasets collected.
To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous quality control process using a cross-validation method. This method involves splitting the document corpora into two main groups, labeled as Group A and Group B. Each group is responsible for annotating page-level and layout-level labels for approximately 1,000 QA pairs. Importantly, there is an overlap of 400 QA pairs between two groups, which serves as a validation checkpoint. The process for quality control works as follows:
  • Overlap Scoring: For 400 overlapping questions, each group’s annotations are scored against the other’s as a form of mutual validation. Specifically, Group A’s annotations are used as a ground truth to score Group B’s work, and conversely, Group B’s annotations serve as the ground truth for scoring Group A.
  • Cross-Evaluation: By cross-evaluating the outcomes from both groups, we are able to assess the consistency of the annotations and identify any discrepancies for improvement. Specifically, we achieved an F1 score of 95.2 for page-level annotations and 87.1 for layout-level annotations, as shown in Table 3.
  • Random Cross-Validation: Apart from the overlapping 400 questions, we randomly cross-validate around 50% of the remaining annotations. In the cases where we have different opinions, we discuss to achieve mutually-agreed annotations. This step ensures the high quality and reliability of MMDocIR.
3.4. Statistics and Analysis
Document Analysis. As shown in Table 3, MMDocIR evaluation set includes 313 long documents averaging 65.1 pages, categorized into ten main domains. Different domains feature distinct distributions of multi-modal information. For instance, research reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich. In contrast, government and legal documents primarily comprise text. Overall, the modality distribution is: Text (60.4%), Image (18.8%), Table (16.7%), and other modalities (4.1%).
Question and Annotation Analysis. MMDocIR encompasses 1,658 questions, 2,107 page labels, and 2,638 layout labels. The modalities required to answer these questions distribute across four categories: Text (44.7%), Image (21.7%), Table (37.4%), and Layout/Meta (11.5%). The “Layout/Meta” category encompasses questions related to layout information and meta-data statistics. Notably, the dataset poses several challenges: 254 questions necessitate cross-modal understanding, 313 questions demand evidence across multiple pages, and 637 questions require reasoning based on multiple layouts. These complexities highlight the need for advanced multi-modal reasoning and contextual understanding.
4. MMDocIR: Training Set
4.1. Document Corpus Collection
The training set corpora are mainly collected from DocVQA benchmarks, adhering to similar selection criteria (see Section 3.1).
  • MP-DocVQA (Tito et al., 2023) contains 47,952 images collected from Industry Documents Library (IDL)  [^5] https://www.industrydocuments.ucsf.edu/. We group the 47,952 document images into separate document files, and obtain 875 long documents (46.8 pages on average) with 15,266 QA pairs.
  • SlideVQA (Tanaka et al., 2023) contains 2,619 slide documents collected from slideshare  [^6] https://www.slideshare.net/ and covering 39 topics. Note that SlideVQA contains only the first 20 pages of each slide decks. In our research, we manually collect the complete slide decks, and obtain 2,011 long documents (averaging 49.3 pages) with 11,066 QA pairs.
  • TAT-DQA (Zhu et al., 2022) consists of 3,067 document pages from financial reports  [^7] https://www.annualreports.com/, dated between 2018 and 2020. Note that neither original documents nor links is provided. We use OCR to extract text in the pages, and use search engine to find relevant documents. After careful tracing and recognition, we identify 163 original documents (averaging 147.3 pages) with 15,814 QA pairs.
  • ArXivQA (Li et al., 2024) comprises 32k figures cropped from academic pages  [^8] https://arxiv.org/. We use the arXiv DOIs provided to collect the academic papers. After careful tracing, recognition, and document length filtering, we identify 1,579 documents averaging 18.4 pages.
  • SciQAG (Wan et al., 2024) consists of 22,728 papers in 24 scientific disciplines, collected from Web of Science (WoS) Core Collection database. We sample 50 documents from each discipline, and manually collect 1,197 papers using the DOIs provided.
  • DUDE (Landeghem et al., 2023) provides 5,019 documents from aggregate websites9991: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/. It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires. We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
  • CUAD (Hendrycks et al., 2021) provides 510 commercial legal contracts, collected from Electronic Data Gathering, Analysis, and Retrieval (EDGAR) [^10] https://www.sec.gov/search-filings. We filter out short documents in CUAD and obtain 274 long documents (29.6 pages on average) with 11,234 QA pairs.
4.2. Label Construction
The page labels can be converted for MP-DocVQA, SlideVQA, and DUDE datasets. Among them, only DUDE provides layout labels.
SciQAG provides only question and answer in texts. We utilize them to reversely obtain the page-level and layout-level labels. Specifically, we first use MinerU to obtain layout-level passage chunks. For each QA pair, we deploy E5 and BGE retrievers to obtain question-passage and answer-passage similarity scores against all extracted passage chunks. If both scores rank within top 3 for a certain passage chunk, we confirm this layout to be the layout-level labels for the given QA pair.
Similarly, ArXivQA provides only cropped images, without document page/layout labels. We first use MinerU to obtain layout-level image. For each cropped image, we calculate the image similarity between all extracted images, and keep only the highest scoring one. Subsequently, we manually examine if the selected image matches the cropped image. In this way, we filter around 20% unmatched images and their QA pairs, resulting 1,579 QA pairs with page-level and layout-level labels.
For TAT-DQA, the layout-level labels are provided for sampled single page. We need to localize the page index of the sampled page. We utilize PDF mapping technique to retrieve the best matched page for every sampled page in this document. Then, we manual examine if the retrieved page is identical to the given page, and correct the labels if there were any errors.
The overall statistics (e.g., document information, modality distribution, domain, etc) of MMDocIR training set are in Table 3.3.
{tabu}
lll—cccccccccc—cc
Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
Report &Indu. Worksh. Paper ure Report book ment Macro Micro
Recall@$k=1$
OCR-text
DPR 21.2 22.1 27.7 23.3 24.4 16.7 21.1 20.7 31.0 15.1 22.3 21.7
ColBERT 43.8 39.8 42.4 39.3 39.2 38.7 46.3 50.6 46.1 17.1 40.3 40.0
BGE 45.5 29.0 41.5 33.6 40.8 32.7 40.0 42.8 36.4 15.1 35.7 35.2
E5 44.2 30.8 39.9 33.2 33.0 32.3 40.4 41.7 38.9 15.8 35.0 34.7
Contriever 39.1 33.3 44.0 34.2 43.9 26.4 40.6 39.4 37.0 15.1 35.3 33.6
GTE 44.6 32.6 45.0 33.2 37.2 31.8 40.0 39.9 35.2 14.5 35.4 34.6
[…]
VLM-text
DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
ColBERT 48.6 42.8 51.1 46.2 36.0 36.8 49.6 60.9 59.5 26.3 45.8 44.9
BGE 48.8 30.9 47.1 40.8 37.6 28.4 43.4 51.9 48.9 28.5 40.6 39.6
E5 48.1 30.0 50.4 39.4 41.1 29.7 40.9 52.8 51.1 24.1 40.8 39.5
Contriever 45.5 31.2 49.8 41.5 39.4 29.4 45.2 55.3 51.1 20.4 40.9 39.7
GTE 46.5 26.3 48.7 38.9 35.9 27.0 46.2 50.1 45.8 24.1 38.9 37.9
[…]
Image
DSEwiki-ss 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
DSEdocmatix 52.3 40.4 56.1 51.7 45.8 43.5 53.8 53.7 58.3 46.7 50.2 50.1
ColPali 56.0 51.8 58.6 55.9 52.0 47.2 57.9 53.9 64.0 32.8 53.0 52.7
DPR-Phi3ours 58.9 50.4 57.4 59.0 57.3 44.6 63.8 50.5 64.4 35.0 54.1 53.7
Col-Phi3ours 56.7 50.4 56.9 61.3 54.8 50.7 60.8 61.3 63.6 54.0 57.0 57.1
Recall@$k=3$
OCR-text
DPR 46.1 40.6 38.9 46.7 43.9 32.4 38.4 37.0 50.0 20.4 39.4 39.8
ColBERT 72.6 59.7 57.8 66.7 60.0 53.7 63.8 68.5 61.4 23.7 58.8 59.5
BGE 69.8 57.7 56.3 58.6 60.7 48.5 57.9 60.9 62.7 20.4 55.4 55.0
E5 66.6 48.7 59.0 58.0 60.9 48.8 63.7 61.4 60.8 20.4 54.8 54.6
Contriever 70.2 55.8 60.4 56.6 62.1 43.0 60.0 56.8 61.4 22.4 54.9 53.6
GTE 69.2 47.0 58.7 59.5 61.8 46.6 65.5 59.1 61.4 19.7 54.9 54.7
[…]
VLM-text
DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
ColBERT 70.1 64.4 70.3 72.3 59.1 55.3 71.1 81.3 70.8 34.3 64.9 64.8
BGE 71.5 48.2 68.8 65.7 56.2 46.5 66.1 69.9 72.0 32.1 59.7 59.6
E5 68.4 45.7 68.1 63.7 60.1 44.0 69.3 72.3 78.8 32.8 60.3 59.3
Contriever 69.4 55.3 68.3 64.9 56.9 46.2 69.9 71.1 72.0 32.1 60.6 59.7
GTE 71.1 44.5 67.2 64.4 54.3 43.0 70.6 71.9 68.2 31.4 58.7 58.3
[…]
Image
DSEwiki-ss 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
DSEdocmatix 75.4 67.5 73.3 80.0 66.3 61.6 72.8 76.4 82.6 57.7 71.4 71.8
ColPali 77.6 71.8 79.4 83.4 72.6 66.1 80.0 80.4 86.4 49.6 74.7 75.0
DPR-Phi3ours 80.3 66.5 77.6 83.9 71.9 63.8 79.8 71.4 84.5 55.5 73.5 74.3
Col-Phi3ours 80.2 74.1 77.4 84.8 69.1 67.7 78.7 79.5 81.8 69.3 76.3 76.8
Recall@$k=5$
OCR-text
DPR 59.5 55.8 43.4 59.1 56.2 41.2 50.7 45.5 56.0 23.0 49.0 49.4
ColBERT 78.4 71.1 63.3 75.2 68.8 60.5 72.0 72.7 67.5 30.3 66.0 66.5
BGE 79.3 65.9 62.1 69.7 69.8 56.5 68.0 62.8 66.3 26.3 62.7 62.9
E5 79.3 62.4 67.0 70.3 71.8 57.6 72.5 67.1 67.5 25.7 64.1 64.2
Contriever 79.9 62.7 64.7 71.7 71.1 48.8 72.4 65.8 67.5 26.3 63.1 62.5
GTE 78.3 61.9 67.3 72.3 68.7 55.2 72.6 64.0 67.5 24.3 63.2 63.5
[…]
VLM-text
DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
ColBERT 78.8 74.0 78.7 82.3 66.1 60.8 77.0 88.5 78.0 38.7 72.3 72.3
BGE 79.5 65.8 71.3 76.8 62.4 56.0 77.2 77.4 79.5 38.0 68.4 68.5
E5 76.9 64.2 75.3 74.4 67.4 52.0 78.5 78.6 82.6 40.9 69.1 67.9
Contriever 77.2 67.1 76.7 75.2 65.1 53.7 75.4 79.2 83.3 39.4 69.2 68.3
GTE 77.4 62.6 74.7 75.8 62.0 51.8 77.8 80.0 75.0 39.4 67.6 67.2
[…]
Image
DSEwiki-ss 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
DSEdocmatix 82.1 77.2 79.6 87.8 73.9 72.4 81.7 83.1 89.4 67.9 79.5 80.1
ColPali 84.6 79.3 82.3 89.0 79.8 72.1 86.7 84.9 92.4 56.9 80.8 81.0
DPR-Phi3ours 86.9 76.2 85.3 91.9 80.0 71.2 87.1 79.5 92.0 61.3 81.1 81.8
Col-Phi3ours 86.3 78.8 81.2 92.4 79.0 73.8 85.3 85.1 87.1 73.0 82.2 83.0
Table 5. Main results for page-level retrieval. “OCR-text” and “VLM-text” refer to converting multi-modal content in the document page using OCR and VLM respectively. “Image” refers to processing document page as screenshot image.
5. Model Training: DPR-Phi3&Col-Phi3
To evaluate the effectiveness of the MMDocIR training set, we train two visual retrievers based on Phi3-Vision (Abdin et al., 2024). Phi3-Vision ($\mathbf{M_{phi3v}}$) reuses the image tokenizer from clip-vit-large [^11] ViT-Large: https://huggingface.co/openai/clip-vit-large-patch14-336 ($\mathbf{M_{vit}}$). It can deal with high-resolution images by cropping them into sub-images, where each sub-image has $336\times 336$ pixels.
Document/Query Encoding
DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR (Karpukhin et al., 2020)) and a list of token-level embeddings (following ColBERT (Khattab and Zaharia, 2020)), respectively. Specifically, we follow Ma et al. (2024a) to concatenate document image with a text prompt: “¡s¿¡d¿ What is shown in this image?¡/s¿”. Here, the ¡d¿ token is a special placeholder token and is replaced by the sequence of patch latent embeddings from the vision encoder. We consider only text queries and use text prompt: “¡s¿ query: ¡q¿ ¡/s¿”. Similarly, the placeholder ¡q¿ token is replaced by input query.
(1) $\begin{split}\mathrm{E_{d}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\\ &\mathrm{E_{q}^{dpr}}=\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\end{split}$
where the end-of-sequence token ¡/s¿ from the last hidden state ($D_{1}=3072$) of $\mathbf{M_{phi3v}}$ is used to represent $\mathrm{E_{d}^{dpr}}$ and $\mathrm{E_{q}^{dpr}}$.
(2) $\begin{split}\mathrm{E_{d}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)},\in\mathbb{R}^{N_{d}\times D_{2}}\\ &\mathrm{E_{q}^{col}}=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)},\in\mathbb{R}^{N_{q}\times D_{2}}\end{split}$
where $\mathbf{M_{proj}}$ is projection layer to map the last hidden states of $\mathbf{M_{phi3v}}$ into reduced dimension $D_{2}=128$. $N_{d}\approx 2500$ for a typical high-resolution page and $N_{q}$ is the number of query tokens.
Query-Doc Similarity
The similarity between the query and the document is computed as follows:
(3) $\text{Sim}(q,d)_{dpr}=\frac{\langle\mathrm{E_{q}^{dpr}}|\mathrm{E_{d}^{dpr}}\rangle}{\left\|\mathrm{E_{q}^{dpr}}\right\|\cdot\left\|\mathrm{E_{d}^{dpr}}\right\|}$
[…]
(4) $\text{Sim}(q,d)_{col}=\sum_{i\in[1,N_{q}]}\max_{j\in[1,N_{d}]}\langle\mathrm{E_{q}^{col}}^{(i)}|\mathrm{E_{d}^{col}}^{(j)}\rangle$
[…]
Contrastive Loss
Given the query $q$, we have the positive document $d^{+}$ and a set of negative documents $d^{-}$ including hard negatives and in-batch negatives. The hard negatives are negative pages within the document with highest $Sim(q,d^{-})$ scored by ColPali (Faysse et al., 2024) retriever. We calculate the loss as:
(5) ($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$
where DPR-Phi3 is trained on the InfoNCE loss, and the temperature parameter $\tau=0.02$ in our experiments.
(6) ($\mathcal{L}^{col}_{(q,d^{+},d^{-})}=\log\Big{(}1+\exp\big{(}\max_{d_{i}\in d^{-}}(\text{Sim}^{col}_{(q,d_{i})})-\text{Sim}^{col}_{(q,d^{+})}\big{)}\Big{)}$
where Col-Phi3 is trained via the softplus loss based on the positive scores w.r.t. to the maximal negative scores.
6. Experiment
{tabu}
llll—cccccccccc—cc
Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
Report &Indu. Worksh. Paper ure Report book ment Macro Micro
Recall@$k=1$
Textual Retrieval
OCR-text
DPR 3.4 7.2 1.2 11.3 3.0 9.8 8.2 24.7 30.9 26.3 12.6 12.4
ColBERT 5.0 8.8 4.7 16.4 2.0 13.2 4.6 50.8 47.7 45.3 19.8 19.1
BGE 7.0 10.9 3.7 14.3 2.3 16.3 8.4 46.1 45.5 35.8 19.0 18.5
E5 6.3 6.0 2.6 14.0 3.5 14.4 6.1 44.7 45.4 40.5 18.4 17.9
Contriever 6.7 7.0 3.8 14.3 4.1 13.3 8.3 43.6 43.9 42.9 18.8 18.1
[…]
VLM-text
DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
ColBERT 22.0 14.9 28.0 28.3 17.9 29.7 21.1 52.6 54.5 44.5 31.3 31.4
BGE 19.2 15.2 24.6 28.7 12.8 27.6 19.7 47.0 52.3 35.8 28.3 29.0
E5 15.9 8.8 27.7 24.3 14.6 21.8 14.7 45.6 53.0 40.5 26.7 26.4
Contriever 23.4 7.5 28.2 26.8 17.1 25.7 16.1 43.6 51.5 42.9 28.3 28.9
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
DSEdocmatix 19.9 11.4 31.5 30.1 17.8 30.0 20.8 46.5 39.4 31.4 27.9 29.1
ColPali 22.5 21.3 36.6 30.9 26.8 32.1 19.3 52.5 51.8 33.6 32.7 32.5
DPR-Phi3ours 21.1 22.1 36.8 35.2 25.6 28.7 24.1 38.3 35.4 27.4 29.5 30.2
Col-Phi3ours 22.6 22.0 37.5 34.9 28.9 30.3 22.7 50.2 45.1 26.3 31.1 31.6
[…]
Hybrid
DSEwiki-ss 14.0 10.4 29.8 18.0 13.7 20.4 13.5 46.0 45.1 34.7 24.6 23.4
DSEdocmatix 18.2 11.6 32.7 24.0 17.7 27.2 16.7 48.1 45.5 33.0 27.5 27.4
ColPali 17.7 12.3 30.0 18.4 19.0 25.5 20.6 49.7 51.2 40.9 28.5 27.1
DPR-Phi3ours 28.3 11.1 35.5 39.2 29.3 25.2 27.4 37.0 32.9 22.2 28.8 30.5
Col-Phi3ours 26.4 12.6 33.7 37.5 30.1 27.9 24.6 46.2 47.4 21.9 30.8 32.0
Recall@$k=5$
Textual Retrieval
OCR-text
DPR 7.3 12.5 5.6 24.0 8.9 16.9 13.6 47.2 50.2 51.1 23.7 23.5
ColBERT 10.9 23.8 10.2 32.2 6.8 25.5 17.0 78.7 63.4 63.2 33.2 32.6
BGE 11.9 20.3 13.6 30.0 11.7 27.7 18.8 68.5 65.4 59.1 32.7 32.2
E5 12.8 16.2 8.9 31.9 10.9 23.6 19.9 76.1 68.8 63.9 33.3 32.7
Contriever 11.9 17.9 11.9 28.8 9.3 24.6 18.1 64.3 68.0 62.7 31.7 31.2
[…]
VLM-text
DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
ColBERT 41.8 37.7 53.7 61.8 35.1 52.4 46.1 83.2 70.1 62.5 54.4 56.0
BGE 41.0 28.1 52.7 59.2 36.7 46.0 50.7 72.0 71.5 59.9 51.8 53.2
E5 35.4 28.1 51.7 58.5 33.2 41.2 40.2 79.7 77.9 64.6 51.1 51.8
Contriever 40.2 29.2 54.1 57.9 36.8 47.1 44.6 68.8 76.2 61.9 51.7 53.0
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
DSEdocmatix 39.6 36.2 53.9 57.5 33.7 52.5 42.8 69.4 63.1 48.9 49.8 51.9
ColPali 40.7 45.9 54.9 58.5 42.6 51.2 45.7 76.8 74.5 48.9 54.0 54.3
DPR-Phi3ours 45.5 37.7 57.0 62.9 41.4 51.1 45.5 65.1 60.8 49.3 51.6 53.9
Col-Phi3ours 46.4 38.2 53.1 61.8 45.0 54.6 45.7 68.8 65.7 43.8 52.3 54.5
[…]
Hybrid
DSEwiki-ss 31.8 29.5 51.1 43.0 34.5 39.3 38.3 71.3 71.3 57.1 46.7 45.6
DSEdocmatix 37.3 26.7 48.2 49.7 34.5 48.6 41.0 72.2 69.4 54.2 48.2 49.3
ColPali 40.1 38.3 55.2 49.2 42.7 47.6 40.8 78.6 68.7 61.0 52.2 51.4
DPR-Phi3ours 54.2 27.4 53.8 64.9 36.6 45.3 49.6 69.5 63.8 53.2 51.8 54.4
Col-Phi3ours 50.9 25.5 49.1 66.2 40.3 48.1 48.3 62.3 60.6 48.5 50.0 53.5
Recall@$k=10$
Textual Retrieval
OCR-text
DPR 10.5 21.1 8.8 32.0 14.9 19.6 17.0 59.7 56.4 58.9 29.9 29.3
ColBERT 14.1 31.1 13.1 38.4 9.1 31.0 22.4 83.9 67.9 65.4 37.6 37.4
BGE 15.7 24.3 15.9 35.9 17.9 31.6 25.3 76.3 73.4 62.1 37.8 37.3
E5 16.9 24.5 13.8 40.1 15.8 26.7 24.7 83.5 77.9 66.1 39.0 38.3
Contriever 15.1 25.7 14.2 36.8 15.9 27.9 24.2 76.8 71.8 64.9 37.3 36.6
[…]
VLM-text
DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
ColBERT 51.0 48.7 60.6 69.8 43.9 61.6 53.7 88.4 74.8 66.4 61.9 63.7
BGE 51.1 38.7 62.1 71.5 41.9 55.6 58.7 80.8 78.7 63.5 60.3 62.4
E5 45.3 38.6 62.0 70.5 45.6 50.0 55.3 87.1 82.4 66.8 60.4 61.2
Contriever 49.9 41.3 62.0 70.5 44.8 56.5 54.5 81.3 78.0 64.9 60.4 62.2
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
DSEdocmatix 53.7 43.3 59.6 66.5 44.7 59.1 50.3 75.4 69.2 53.7 57.5 59.9
ColPali 53.6 54.1 64.4 69.5 48.8 60.7 54.0 81.9 82.5 50.4 62.0 63.2
DPR-Phi3ours 58.1 49.1 67.0 74.7 48.4 57.9 57.8 68.7 66.2 54.4 60.2 62.8
Col-Phi3ours 57.7 50.5 66.6 72.3 50.7 59.3 53.6 68.5 74.8 57.5 61.1 63.3
[…]
Hybrid
DSEwiki-ss 44.1 34.3 57.6 56.3 42.6 50.7 48.6 81.1 79.1 63.7 55.8 56.0
DSEdocmatix 49.9 37.3 57.3 61.3 45.9 57.9 50.1 77.9 74.9 61.5 57.4 58.9
ColPali 52.1 46.4 65.0 64.4 50.7 53.8 51.0 82.7 71.4 62.5 60.0 60.2
DPR-Phi3ours 65.2 33.7 60.3 72.9 42.4 54.1 52.9 79.1 72.5 65.6 59.9 62.9
Col-Phi3ours 59.1 38.7 57.0 77.7 43.9 57.7 51.7 72.4 68.0 60.7 58.7 62.8
Table 6. Main results for layout-level retrieval. “OCR-text” and “VLM-text” refer to converting multi-modal layouts using OCR and VLM respectively. “Pure-Image” and “Hybrid” refer to reading textual layouts in image and text format respectively.
6.1. Evaluation Metric
The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores. Recall@$k$ is defined as the proportion of the ground truth page/layout evidence that is successfully retrieved. Note for layout matching, we calculate recall based on the overlaps of bounding boxes between retrieved layouts and gold layouts.
6.2. Baseline Models and Setting
We evaluate 6 state-of-the-art text retrievers: namely DPR (Karpukhin et al., 2020), ColBERT (Khattab and Zaharia, 2020), BGE (Xiao et al., 2023), E5 (Wang et al., 2022), Contriever (Izacard et al., 2022), and GTE (Li et al., 2023) (refer to Section 7.1). Meanwhile, we evaluate 5 VLM-based retrievers: 3 off-the-shelf models named DSEwiki-ss (Ma et al., 2024a), DSEdocmatix (Ma et al., 2024a), and ColPali (Faysse et al., 2024) (refer to Section 7.2), and 2 our trained models (see section 5). All retrievers are adapted into dual-task setting:
  • Page Retrieval: For textual retrievers, we use MinerU (Wang et al., 2024a) to extract the layout of all document pages. Non-textual layouts are converted into ‘OCR-text” and “VLM-text” by using Tesseract OCR  [^12] https://github.com/tesseract-ocr/tesseract and GPT-4o respectively. The document page is represent by concatenating the text of all layouts within it. For visual retrievers, we directly utilize document page screenshots.
  • Layout Retrieval: Textual retrievers process non-textual layouts using “OCR-text” or “VLM-text”. Visual retrievers process textual layouts in either “Pure Image” (using cropped image of textual area as input) or “Hybrid” (using its original text, as VLM can directly encode text as well).
6.3. Main Results for Page-level Retrieval
For page-level retrieval (Table 4.2), our key findings are as follows:
  • Superiority of Visual Retrievers: Visual retrievers consistently outperform text retrievers across various domains and retrieval metrics (e.g., Top $k=1,3,5$). This highlights the significance of leveraging document screenshots to capture multi-modal information, which is often lost when relying solely on OCR-text.
  • Effectiveness of MMDocIR: Among the visual retrievers, DPR-Phi3ours and Col-Phi3ours trained on theMMDocIR train set demonstrate superior performance. It suggests dataset quality to effectively enhances the retrieval ability.
  • Effectiveness of VLM-Text: Although VLM-text approaches underperform visual retrievers, they achieve much better performance than the OCR-text methods. This indicates benefits of using GPT-4o to preserve visual cues in text.
  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), achieve more advantageous results in Recall@1 and marginal performance increase in Recall@5/10. Yet, the storage overhead of token-level embeddings can be 10 times more than page-level embeddings.
  • Top 5 Coverage: Retrieving top 5 pages provides substantial coverage, ensuring that relevant information is captured.
6.4. Main Results for Layout-level Retrieval
For layout-level retrieval (Table 6), our key findings are as follows:
  • Superiority of Visual Retrievers: Visual retrievers exhibit substantial performance advantages over text retrievers utilizing OCR-text. This highlights the limitations of OCR in capturing the multi-modal nature of documents, where visual context can significantly enhance retrieval accuracy.
  • Effectiveness of VLM-Text: Interestingly, VLM-text approaches achieve comparable performance as visual retrievers. This indicates strong image description capabilities of state-of-the-art VLM, which can offer significant benefits to textual retrievers in understanding multi-modal information.
  • Comparison of Hybrid vs. Pure Image Sequences: Visual retrievers relying on hybrid image-text sequences generally perform less effectively than those utilizing pure image sequences. This suggests that current VLMs may have stronger capabilities in modeling images than text within the multi-modal framework.
  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), do not achieve advantageous results in layout retrieval.
  • Top 10 Coverage: Retrieving top 10 layouts does not provide comprehensive coverage of the ground truth layouts for document retrieval tasks. It suggests the challenge of layout retrieval.
6.5. Analysis of OCR and VLM Text
Interestingly, text retrievers leveraging VLM-text can significantly outperform OCR-text, in both page and layout retrieval. OCR-text which is the the raw text extracted via OCR tools, and as experimental results suggest, it is not suitable for multi-modal retrieval. In comparison, VLM-text can facilitate multi-modal retrieval, suggesting that VLMs can largely preserve rich multi-modal information. The averaged word length and distribution of OCR and VLM text obtained from the tables and figures in MMDocIR in shown in Figure 3. Observe that the VLM-text is much longer than OCR-text. Specifically, the length of VLM-text is 1.5 and 3.8 times more than that of OCR-text in table and figure. Tables tend to contain numerical or structured layouts, which can be well recorded by text. Figures contain mostly graphical elaboration and visual cues. Hence, making OCR tools struggle in getting textual information from figures. Although VLM-text can offer much comprehensive text information than OCR-text, it comes at the cost of higher computational overhead and longer latency.
(a) Avg word length
(b) Distribution density of word length
Figure 3. Avg length and distribution of OCR and VLM text.
7. Related Work
7.1. Text-Centric Document Retrieval
Document indexing.
The process of indexing a multi-modal document involves multiple steps, including Document Parsing (Chao and Fan, 2004; Wang et al., 2024a), Optical Character Recognition (OCR) (Chaudhuri et al., 2017; Borovikov, 2014; Mori et al., 1999), Layout Detection (Sassioui et al., 2023; Xu et al., 2020, 2021), Chunking (Chen et al., 2024a; Raina and Gales, 2024; Dong et al., 2024), and Image Captioning (You et al., 2016; Aneja et al., 2018). These steps are time-consuming and can introduce errors that impact the overall retrieval performance.
Text retrieval.
Current text retrieval are primarily classified into sparse and dense retrieval. For two widely-used sparse retrievers: TF-IDF (Salton et al., 1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 (Robertson et al., 1994) introduces nonlinear word frequency saturation and length normalization. Dense retrievers encode content into vector representations. DPR (Karpukhin et al., 2020) is the pioneering work of dense vector representations for QA tasks. Similarly, ColBERT (Khattab and Zaharia, 2020) introduces an efficient question-document interaction model with late fine-grained term matching. Contriever (Izacard et al., 2022) leverages contrastive learning to improve content dense encoding. E5 (Wang et al., 2022) and BGE (Xiao et al., 2023) propose novel training and data preparation techniques to enhance retrieval performance. Moreover, GTE (Li et al., 2023) integrates graph-based techniques to enhance dense embedding. Most systems focus on text-based representations, neglecting the valuable visual information present in documents.
7.2. Vision-Driven Document Retrieval
Vision Language Models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) have gained popularity for their ability to understand and generate text based on combined text and visual inputs. This advancement has led to the development of cutting-edge visual-driven retrievers, such as ColPali (Faysse et al., 2024) and DSE (Ma et al., 2024a). These models specifically leverage PaliGemma (Beyer et al., 2024) and Phi3-V (Abdin et al., 2024) to directly encode document page screenshots for multi-modal document retrieval. ColPali adopts similar question-document interaction as ColBERT, and represent each document page in token-level embeddings. By contrast, DSE is similar to DPR to encode each page with a single dense embedding. Visual Retrievers can directly model useful visual information, allowing for the direct utilization of multi-modal content without first converting it into text. Despite these advancements, visual retrievers face challenges, particularly in dealing with text details when document page resolutions are high. The high resolution of document pages substantially increases the computational cost and complexity of the embedding process, which may hinder the model’s performance.
7.3. Multi-modal Document Retrieval Datasets
As described in Section 1 and Table 1, there is a notable lack of a robust benchmark for multi-modal document retrieval. DocCVQA (Tito et al., 2021) is the first multi-modal document retrieval-answering task, which extracts information from a document image collection and then provides the answer. However, DocCVQA provides only 20 questions, and the corpora is limited to finance domain specifically in political disclosure. PDF-MVQA (Ding et al., 2024) is tailored for multi-modal retrieval in research articles. Note that PDF-MVQA is annotated by Chat-GPT (GPT-3.5-turbo) but not experts, and its corpora contains only biomedical articles. SciMMIR (Wu et al., 2024) proposes multi-modal retrieval in scientific research papers. However, it provides only the image-caption pairs, not the natural user query with the targeting document pages. Ma et al. (2024a) introduce two relevant datasets, namely Wiki-SS and DocMatix-IR. Wiki-SS is constructed based on Natural Questions (Kwiatkowski et al., 2019) where the evidence passage is the screenshot wikipedia webpage. Natural Questions are mainly designed for text rather than multi-modal retrieval. We also notice that the screenshot may not include the ground-truth evidence as only the front-page is screenshotted. DocMatix-IR is constructed from the largest document understanding dataset DocMatix (Laurençon et al., 2024) via filtering and hard negative mining. However, the VQA questions in DocMatix is constructed by Phi-3-small (Abdin et al., 2024) rather than experts, and these VQA questions are not de-contextualized for retrieval.
ViDoRe (Faysse et al., 2024) is the most relevant benchmark to MMDocIR. It integrates multiple DocVQA datasets (Mathew et al., 2021, 2022; Zhu et al., 2022; Li et al., 2024), and provides new documents in scientific, medical, administrative, and environment domains. Upon thorough examination of the 1,180 questions within the ViDoRe dataset, we find that more than 80% of these questions present drawbacks in terms of their complexity and context-specific nature. Typical examples include “What’s the title of the picture” and “what’s the author’s name”. They tend to be either overly simplistic or excessively dependent on a specific context which does not necessarily reflect the broader, more complex needs of effective IR. Additionally, a significant limitation of the ViDoRe dataset is its reliance on sampled document pages or cropped images rather than providing the complete document corpora.
8. Conclusion
In conclusion, multi-modal document retrieval presents a complex challenge that necessitates the integration of diverse data modalities beyond plain text. Our contributions address this challenge by introducing the MMDocIR benchmark, which includes innovative dual-task retrieval functionalities targeting both page-level and layout-level specifics of documents. The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation purposes. Our comprehensive evaluation reveals that visual-driven retrievers significantly outperform text-driven ones, underscoring the critical importance of incorporating visual content in enhancing the effectiveness of document retrieval systems. Future work can expand on optimizing these retrieval algorithms to improve the accuracy and efficiency of multi-modal document retrieval systems.